# Use as Many Surrogates as You Want: Selective Ensemble Attack to Unleash Transferability without Sacrificing Resource Efficiency

### v1

@@ 4.2 SEA in Various Transfer Scenarios
Our SEA gives the potential to use many models without increasing the resource costs. To explore the potential of utilizing many models in SEA, we assume easy access to 20 pre-trained models from the PyTorch library. As can be seen from Figure 3, when varying the within-iteration model $m$ from 1 to 20, our SEA(20,$m$) consistently surpasses Ens(20,$m$) in all cases. The performance gap becomes gradually larger as $m$ decreases. This trend indicates that the superiority of our SEA stands out especially when the resource is restricted in practice. In particular, when the common $m=4$ is used, our SEA(20,4) outperforms Ens(20,4) by 8.5% on average. More detailed results for (20,4) on various target models are reported in Table 3, and results for (4,3), a more commonly used ensemble setting are in Table 12 of Appendix E.
@@
We have the following observations. First, SEA performs better on all diverse (black-box) target models including CNNs, ViTs, and Defenses. This is because SEA involves more models across iterations (but without consuming more resources). Second, Ens(4,3) and SEA(4,3) consume very close resources, because they have the same within-iteration models. Third, Table 3 shows different transfer baselines consume different costs, and more costs generally lead to more effective attacks.
@@
SEA vs. advanced ensemble strategies. We further compare our SEA to more complex, advanced model ensemble strategies, such as SVRE, AdaEA, SMER, and CWA. These strategies optimize the ensemble attack from different perspectives (see descriptions in Section 2), beyond equally averaging the logit or loss values of surrogate models, and follow Ens to use identical models across iterations. As can be seen from Table 4, our SEA still achieves the best transferability. For resource efficiency, the simplicity of the random model selection makes SEA consume substantially less computational time and memory. We also find that combining the above complex ensemble strategies with our SEA can further improve transferability. See detailed results in Table 14 of Appendix F.
@@
Attacking real-world systems. We further compare Ens and SEA in attacking real-world systems: the commercial vision APIs, Google and Baidu Cloud Vision, and the large vision-language models (LVLM), Qwen and ChatGLM. Notably, given an (original or adversarial) image, we prompt the LVLMs using "Output a single-word label in English for this image." An adversarial image is considered successful when its output label differs from that of the original image. Table 6 and 6 show that SEA achieves better results than Ens on both APIs and LVLMs, with the MI transfer baseline. Figure 4 further shows successful SEA image examples on both APIs. Additional visualizations, also on two LVLMs, are provided in the Appendix I, where for LVLMs, we also try a finer-grained testing with the prompt "Describe the image no more than 50 words".
@@
Targeted attacks. We further compare SEA and Ens under more challenging targeted attacks. Specifically, we follow the common practice [41, 48], using 100 iterations and a step size of 2, with DI-TI-MI as the baseline. The results in Table 7 demonstrate that SEA consistently outperforms Ens. For instance, SEA(4,2) achieves an 8.3% improvement in average attack success rate while reducing resource consumption compared to Ens(4,3), highlighting the superiority of our method.
@@ 4.3 Ablation Studies
Base Sampling strategy ASR (%) Time (min) Memory (GB)
@@
Diversity of surrogate models. To validate the general effectiveness of our SEA, we consider ensemble settings with similar or diverse surrogate models. Specifically, we adopt four ensemble models from the resnet, vgg, densenet, or dpn family, as well as their combination, diverse. As can be seen from Table 8, our SEA(4,3) consistently achieves higher transferability than Ens(4,3) across different model combinations. In addition, using a more diverse combination of surrogate models yields generally higher transferability for both Ens and SEA. We also validate challenging cases with the ensemble of CNNs and ViTs in Appendix H, showing that even with such highly diverse model architectures, there are no opposite effects [2].
Model selection strategy. Selecting a subset of models from all models in each iteration results in an enormous search space. For example, when selecting 2 models from 20, there are $190^{10}$ possibilities over 10 iterations. Here we compare whether different selection strategies may have an impact on SEA. Specifically, we compare random (with or without replacement across iterations) selection and an optimization-based selection [51, 26], which updates the probability of each model being selected through differential augmentation. As can be seen from Table 9, random sampling without replacement yields the best results. This is because it avoids selecting the same models as those in previous iterations, leading to higher cross-iteration model diversity. For example, for SEA(20,2), random selection without replacement for 10 iterations yields the number of cross-iteration models $n=s=20$. The stability of random sampling is analyzed in Table 10 of Appendix C.
@@ 5 Conclusion and Outlook
Appendix A Limitations and Broader Impacts
In this work, we have explored and addressed the trade-off between transferability and resource efficiency. Another important property of (transfer-based) adversarial examples is the stealthiness of modified images, which has been simply restricted using an $L_{\infty}$ bound following existing work. In the future, we would explore whether it is possible to better trade off transferability and stealthiness.
Like other adversarial attack methods, if the proposed SEA method is maliciously used in practice, it may lead to security issues in deep learning models, representing one of its potential negative societal impacts. On the bright side, by leveraging the attack strength of adversarial examples, SEA can serve as a useful tool for evaluating and selecting the most robust models in safety-critical applications. In particular, the efficiency of our SEA would help speed up the evaluation and testing process. In addition, our study can provide researchers and practitioners constructive guidance for developing adversarial defenses.
Appendix B Analysis of the Expectation of $n$
@@
Appendix C Stability Testing of Random Sampling in SEA
@@
We conduct an in-depth analysis of the stability of random sampling. Specifically, we repeat the experiments of SEA and Ens 10 times using the MI baseline. The experimental results in Table 10 show that Ens(4,3) and SEA(4,3) exhibit comparable and low variability, indicating that the random sampling strategy offers good stability.
Appendix D Model Loading Time
@@
Appendix E Additional Experiments in Varied Settings
@@
Baseline Attack CNNs ViTs Defense Avg. Time (min) Memory (GB)
@@
Figure 8: Attack success rates (%) of transferring from an ensemble of $m$ models to eight different target models. TI is the transfer baseline.
@@
Surrogates Attack CNNs ViTs Defense Avg. Time (min) Memory (GB)
V19 D121 ViT-B Swin-B ATens4 ATens HGD RS
Ens(4,4) 96.5 99.1 61.8 48.7 84.0 76.7 90.4 82.4 80.0 1.9 7.2
Ens(4,3) 95.4 98.2 57.6 40.1 78.2 70.9 85.9 78.6 75.6 1.6 6.8
SEA(4,3) 96.9 98.8 61.1 46.8 83.2 76.6 90.9 82.3 79.6 1.5 6.8
resnet SEA(4,2) 96.1 98.8 59.7 43.7 81.4 74.0 88.7 82.3 78.1 1.2 6.4
Ens(4,4) 99.9 95.6 52.9 43.5 70.8 62.5 78.7 72.5 72.1 2.7 10.6
Ens(4,3) 99.8 94.4 49.4 38.7 67.8 58.0 72.2 68.8 68.6 2.1 10.1
SEA(4,3) 99.9 95.8 53.0 42.4 70.4 60.7 76.5 71.8 71.3 2.1 8.9
vgg SEA(4,2) 99.9 94.3 52.3 40.6 69.2 58.7 75.0 71.7 70.2 1.1 7.5
Ens(4,4) 94.7 99.9 66.0 53.5 85.1 81.9 92.5 81.4 81.9 3.2 16.3
Ens(4,3) 92.6 99.9 62.6 47.8 81.8 75.3 89.8 77.8 78.5 2.5 13.7
SEA(4,3) 94.3 99.9 64.7 52.3 84.8 80.8 91.8 81.6 81.3 2.5 13.6
densenet SEA(4,2) 94.0 99.9 63.5 48.2 83.6 79.0 90.4 81.1 80.0 1.9 10.7
Ens(4,4) 85.7 95.5 64.0 52.7 85.3 85.2 90.9 74.7 79.3 6.6 22.0
Ens(4,3) 83.2 94.5 58.1 46.4 80.1 77.1 86.0 72.1 74.7 5.2 19.9
SEA(4,3) 85.8 95.4 63.5 51.6 85.8 83.6 90.2 74.9 78.9 5.1 20.1
dpn SEA(4,2) 85.0 95.1 61.4 51.3 84.1 81.6 88.4 74.8 77.7 3.7 15.4
Ens(4,4) 97.3 98.3 68.2 57.9 87.2 83.5 92.8 82.5 83.5 4.5 17.8
Ens(4,3) 94.8 98.1 60.8 49.1 81.6 77.8 90.2 78.3 78.8 3.6 16.5
SEA(4,3) 97.2 98.6 66.3 56.4 86.5 83.1 92.2 82.1 82.8 3.5 15.9
diverse SEA(4,2) 97.1 98.5 63.7 54.7 85.4 81.9 91.0 80.8 81.6 2.6 14.1
@@
Base Ensemble strategy Attack CNNs ViTs Defense Avg. Time (min) Memory (GB)
@@
We provide additional experimental results to demonstrate the validity of the conclusions presented in the main text. These experiments cover varied number of surrogate models, surrogate model architectures, and ensemble strategies. For experiments that have been reported with the MI baseline in the main text, we additionally evaluate TI, a popular input transformation-based transferable attack. We also add another popular surrogate model refinement attack baseline [23, 42, 49], ghost [23]. Ghost randomly drops out neurons in the surrogate model.
@@
Appendix F SEA Integrated with Advanced Ensemble Strategies
In the main text, we have compared our SEA to advanced model ensemble strategies: SVRE, AdaEA, SMER, and CWA. Here we further show that SEA can integrate them for further improvement of transferability. As can be seen from Table 14, our SEA consistently outperforms Ens by a large margin. For example, our SEA boosts the transferability of the state-of-the-art method, CWA, by 6.0% on average.
@@ Appendix F SEA Integrated with Advanced Ensemble Strategies
Appendix G The Impact of Batchsize on Resource Efficiency
@@
Appendix H Ensemble of ViTs and CNNs
@@
We further validate the effectiveness of SEA under a more challenging ensemble setting involving both CNNs and ViTs, as opposite effects may occur when highly diverse models are selected [2]. Specifically, Ens ensembles three CNNs, whereas SEA adds a model with a different architecture to the ensemble. The results in Table 15 show that even when the ensemble models have diverse architectures (CNN and ViT), SEA still outperforms Ens on all target models. More specifically, incorporating an additional CNN (ViT) into the ensemble enhances the attack performance against CNN (ViT) target models.
Appendix I Visualizations of Adversarial Images on APIs and VLMs
In Section 4.2, the quantitative results on APIs and LVLMs (presented in Tables 6 and 6) have demonstrated that SEA outperforms Ens. Here, we provide visualizations of the original images and the adversarial images crafted by SEA, as shown in Figures 10, 11, 12, 13, and 14. It can be observed that even ensembling small image classification models can effectively attack APIs and LVLMs, posing significant challenges in the development of robust vision APIs and LVLMs.
@@ Appendix I Visualizations of Adversarial Images on APIs and VLMs
Figure 11: Original image and SEA adversarial image on Qwen with the prompt "Output a single-word label in English for this image".
Figure 12: Original image and SEA adversarial image on ChatGLM with the prompt "Output a single-word label in English for this image".
Figure 13: Original image and SEA adversarial image on Qwen with the prompt "Describe the image, no more than 50 words".
Figure 14: Original image and SEA adversarial image on ChatGLM with the prompt "Describe the image, no more than 50 words".

### v2

@@ 4.2 SEA in Various Transfer Scenarios
Our SEA gives the potential to use many models without increasing the resource costs. To explore the potential of utilizing many models in SEA, we assume easy access to 20 pre-trained models from the PyTorch library. As can be seen from Figure 3, when varying the within-iteration model $m$ from 1 to 20, our SEA(20,$m$) consistently surpasses Ens(20,$m$) in all cases. The performance gap becomes gradually larger as $m$ decreases. This trend indicates that the superiority of our SEA stands out especially when the resource is restricted in practice. In particular, when the common $m=4$ is used, our SEA(20,4) outperforms Ens(20,4) by 8.5% on average. More detailed results for (20,4) on various target models are reported in Table 3, and results for (4,3), a more commonly used ensemble setting are in Table 11 of Appendix D.
@@
We have the following observations. First, SEA performs better on all diverse (black-box) target models including CNNs, ViTs, and Defenses. This is because SEA involves more models across iterations (but without consuming more resources). Second, Ens(4,3) and SEA(4,3) consume very close resources, because they have the same within-iteration models. Third, Table 11 shows that different transfer baselines consume different costs, and more costs generally lead to more effective attacks.
@@
SEA vs. advanced ensemble strategies. We further compare our SEA to more complex, advanced model ensemble strategies, such as SVRE, AdaEA, SMER, and CWA. These strategies optimize the ensemble attack from different perspectives (see descriptions in Section 2), beyond equally averaging the logit or loss values of surrogate models, and follow Ens to use identical models across iterations. As can be seen from Table 4, our SEA still achieves the best transferability. For resource efficiency, the simplicity of the random model selection makes SEA consume substantially less computational time and memory. We also find that combining the above complex ensemble strategies with our SEA can further improve transferability. See detailed results in Table 13 of Appendix E.
@@
Attacking real-world systems. We further compare Ens and SEA in attacking real-world systems: the commercial vision APIs, Google and Baidu Cloud Vision, and the large vision-language models (LVLM), Qwen and ChatGLM. Notably, given an (original or adversarial) image, we prompt the LVLMs using ”Output a single-word label in English for this image.” An adversarial image is considered successful when its output label differs from that of the original image. Table 5 shows that SEA achieves better results than Ens on both APIs and LVLMs, with the MI transfer baseline. Figure 4 further shows successful SEA image examples on both APIs. Additional visualizations, also on two LVLMs, are provided in the Appendix H, where for LVLMs, we also try a finer-grained testing with the prompt ”Describe the image no more than 50 words”.
@@
Targeted attacks. We further compare SEA and Ens under more challenging targeted attacks. Specifically, we follow the common practice [wei2023enhancing, zhao2021success], using 100 iterations and a step size of 2, with DI-TI-MI as the baseline. The results in Table 6 demonstrate that SEA consistently outperforms Ens. For instance, SEA(20,3) achieves an 32.0% improvement in average attack success rate while reducing resource consumption compared to Ens(20,4), highlighting the superiority of our method.
@@ 4.3 Ablation Studies
Base
Sampling
strategy
ASR
Time
Memory
@@
model scheduling 86.7 48.7 12.8
Diversity of surrogate models. To validate the general effectiveness of our SEA, we consider ensemble settings with similar or diverse surrogate models. Specifically, we adopt four ensemble models from the resnet, vgg, densenet, or dpn family, as well as their combination, diverse. As can be seen from Table 7, our SEA(4,3) consistently achieves higher transferability than Ens(4,3) across different model combinations. In addition, using a more diverse combination of surrogate models yields generally higher transferability for both Ens and SEA. We also validate challenging cases with the ensemble of CNNs and ViTs in Appendix G, showing that even with such highly diverse model architectures, there are no opposite effects [chen2023AdaEA].
@@
Model selection strategy. Selecting a subset of models from all models in each iteration results in an enormous search space. For example, when selecting 2 models from 20, there are $190^{10}$ possibilities over 10 iterations. Here, we compare whether different selection strategies may have an impact on SEA. Specifically, we compare three model selection strategies: 1) our random selection (with or without replacement across iterations); 2) an optimization-based selection [zhu2024learning, liu2021direct], which updates the probability of each model being selected through differential augmentation; 3) a model scheduling approach [hu2022model], which prioritizes models with representative decision boundaries (measured by gradient direction outlier degree) while progressively deprioritizing previously selected or defeated models across attack stages. As can be seen from Table 8, random sampling without replacement yields the best results. This is because it avoids selecting the same models as those in previous iterations, leading to higher cross-iteration model diversity. For example, for SEA(20,2), random selection without replacement for 10 iterations yields the number of cross-iteration models $n=s=20$. The stability of random sampling is analyzed in Table 9 of Appendix B.
@@ 5 Conclusion and Outlook
Appendix A Analysis of the Expectation of $n$
@@
Appendix B Stability Testing of Random Sampling in SEA
@@
We conduct an in-depth analysis of the stability of random sampling. Specifically, we repeat the experiments of SEA and Ens 10 times using the MI baseline. The experimental results in Table 9 show that Ens(4,3) and SEA(4,3) exhibit comparable and low variability, indicating that the random sampling strategy offers good stability. It should be noted that the four-model ensemble employs Incv3 [Szegedy2016inc-v3], Incv4, IncResv2 [Szegedy2017incRes], and Res152 [He2016residual] as surrogate models, consistent with prior works [Lin2020Nesterov, Wang2021Variance]. To ensure a black-box evaluation, Vgg19 (V19) and Densenet121 (D121) are used as CNN test (target) models, while other test models remain unchanged.
Appendix C Model Loading Time
@@
Appendix D Additional Experiments in Varied Settings
@@
Baseline Attack CNNs ViTs Defense Avg.
Time
Memory
@@
Figure 8: Attack success rates (%) of transferring from an ensemble of $k$ models to eight different target models. TI is the transfer baseline.
@@
Surrogates Attack CNNs ViTs Defense Avg.
Time
Memory
Incv4 Res152 ViT-B Swin-B ATens4 ATens HGD RS
Ens(4,4) 88.0 90.8 61.8 48.7 84.0 76.7 90.4 82.4 77.9 1.9 7.2
Ens(4,3) 81.7 86.7 57.6 40.1 78.2 70.9 85.9 78.6 72.5 1.6 6.8
SEA(4,3) 88.0 91.1 61.1 46.8 83.2 76.6 90.9 82.3 77.5 1.5 6.8
resnet SEA(4,2) 86.5 89.5 59.7 43.7 81.4 74.0 88.7 82.3 75.7 1.2 6.4
Ens(4,4) 89.3 77.9 52.9 43.5 70.8 62.5 78.7 72.5 68.5 2.7 10.6
Ens(4,3) 86.7 74.6 49.4 38.7 67.8 58.0 72.2 68.8 64.5 2.1 10.1
SEA(4,3) 89.1 77.0 53.0 42.4 70.4 60.7 76.5 71.8 67.6 2.1 8.9
vgg SEA(4,2) 87.2 76.8 52.3 40.6 69.2 58.7 75.0 71.7 66.4 1.1 7.5
Ens(4,4) 90.7 92.2 66.0 53.5 85.1 81.9 92.5 81.4 80.4 3.2 16.3
Ens(4,3) 86.9 89.1 62.6 47.8 81.8 75.3 89.8 77.8 76.4 2.5 13.7
SEA(4,3) 90.6 91.8 64.7 52.3 84.8 80.8 91.8 81.6 79.8 2.5 13.6
densenet SEA(4,2) 89.5 91.1 63.5 48.2 83.6 79.0 90.4 81.1 78.3 1.9 10.7
Ens(4,4) 88.8 90.0 64.0 52.7 85.3 85.2 90.9 74.7 79.0 6.6 22.0
Ens(4,3) 83.2 86.9 58.1 46.4 80.1 77.1 86.0 72.1 73.7 5.2 19.9
SEA(4,3) 88.4 89.9 63.5 51.6 85.8 83.6 90.2 74.9 78.5 5.1 20.1
dpn SEA(4,2) 87.8 88.8 61.4 51.3 84.1 81.6 88.4 74.8 77.3 3.7 15.4
Ens(4,4) 91.2 92.5 68.2 57.9 87.2 83.5 92.8 82.5 82.0 4.5 17.8
Ens(4,3) 88.3 89.8 60.8 49.1 81.6 77.8 90.2 78.3 77.0 3.6 16.5
SEA(4,3) 91.7 91.8 66.3 56.4 86.5 83.1 92.2 82.1 81.3 3.5 15.9
diverse SEA(4,2) 91.5 91.3 63.7 54.7 85.4 81.9 91.0 80.8 80.0 2.6 14.1
@@
Base
Ensemble
strategy
Attack CNNs ViTs Defense Avg.
Time
Memory
@@
We provide additional experimental results to demonstrate the validity of the conclusions presented in the main text. These experiments cover varied number of surrogate models, surrogate model architectures, and ensemble strategies. The setting of 4 surrogate models is consistent with that of previous works [Lin2020Nesterov, Wang2021Variance]. For experiments that have been reported with the MI baseline in the main text, we additionally evaluate TI, a popular input transformation-based transferable attack. We also add another popular surrogate model refinement attack baseline [li2020learning, wu2020skip, zhao2025revisiting], Ghost [li2020learning]. Ghost randomly drops out neurons in the surrogate model.
@@
Appendix E SEA Integrated with Advanced Ensemble Strategies
In the main text, we have compared our SEA to advanced model ensemble strategies: SVRE, AdaEA, SMER, and CWA. Here we further show that SEA can integrate them for further improvement of transferability. As can be seen from Table 13, our SEA consistently outperforms Ens by a large margin. For example, SEA boosts the transferability of the state-of-the-art method, CWA, by 6.0% on average.
@@ Appendix E SEA Integrated with Advanced Ensemble Strategies
Appendix F The Impact of Batchsize on Resource Efficiency
@@
Appendix G Ensemble of ViTs and CNNs
@@
We further validate the effectiveness of SEA under a more challenging ensemble setting involving both CNNs and ViTs, as opposite effects may occur when highly diverse models are selected [chen2023AdaEA]. Specifically, Ens ensembles three CNNs, whereas SEA adds a model with a different architecture to the ensemble. The results in Table 14 show that even when the ensemble models have diverse architectures (CNN and ViT), SEA still outperforms Ens on all target models. More specifically, incorporating an additional CNN (ViT) into the ensemble enhances the attack performance against CNN (ViT) target models.
Appendix H Visualizations of Adversarial Images on APIs and VLMs
In Section 4.2, the quantitative results on APIs and LVLMs (presented in Tables 5) have demonstrated that SEA outperforms Ens. Here, we provide visualizations of the original images and the adversarial images crafted by SEA, as shown in Figures 10, 11, 12, 13, and 14. It can be observed that even ensembling small image classification models can effectively attack APIs and LVLMs, posing significant challenges in the development of robust vision APIs and LVLMs.
@@ Appendix H Visualizations of Adversarial Images on APIs and VLMs
Figure 11: Original image and SEA adversarial image on Qwen with the prompt ”Output a single-word label in English for this image”.
Figure 12: Original image and SEA adversarial image on ChatGLM with the prompt ”Output a single-word label in English for this image”.
Figure 13: Original image and SEA adversarial image on Qwen with the prompt ”Describe the image, no more than 50 words”.
Figure 14: Original image and SEA adversarial image on ChatGLM with the prompt ”Describe the image, no more than 50 words”.